The sum of two integers is 5. The sum of their cubes is 35. What is the sum of their squares?
Answer by richard1234(7193)

2+3 = 5 and 2^3 + 3^3 = 35

The integers are 2,3. The sum of their squares is 2^2 + 3^2 = 13.